OpenStudy (anonymous):

question.. Which ratio is equivalent to 6 : 7? A. 12 : 13 B. 18 : 48 C. 36 : 49 D. 48 : 56

OpenStudy (anonymous):

@tosin please help xD i guess

OpenStudy (anonymous):

a

Directrix (directrix):

Does this: 18 : 48 simplify to 6:7 ?

OpenStudy (anonymous):

Yeah, I though A because 6 12 18 for B is 3x which is not that big

OpenStudy (cpinder):

simplify all of the ratios and then choose which one comes out to be 6:7.

OpenStudy (cpinder):

A is not correct.

Directrix (directrix):

A is wrong. We have users on this site who throw out answers with no work. Frequently, those answers are wrong as is this one. @MyPhilHillName

OpenStudy (anonymous):

Wait.....OHHH its b OK I GOT IT

Directrix (directrix):

What is the reason? Does 18/48 = 6/7 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

b rather, was a mistake .

OpenStudy (anonymous):

mistype much? xD

OpenStudy (anonymous):

lol, dont bite my head off. i just reviewed it.

OpenStudy (cpinder):

18/48 does not simplify to 6/7. if you divide 18 by 3, you get 6, but if you divide 48 by 3, you get 16. so your ratio is 6/16 which is not 6/7.

OpenStudy (anonymous):

haha, @MyPhilHillName

OpenStudy (cpinder):

how about trying to simplify 48:56 by dividing both numbers by 8

OpenStudy (anonymous):

-_- ima row my boat down the stream

OpenStudy (cpinder):

divide 48 by 8 and then 56 by 8. @MyPhilHillName

OpenStudy (anonymous):

i thought C because 49 divided by 7 is 7 kthxbai

OpenStudy (anonymous):

and 36 divided by 6 is 6

OpenStudy (cpinder):

I'll try to explain this to you better. You have to pick one number to divide both numbers by. So if you divide 49 by 7, you get 7. But, you have to divide 36 by 7 as well. 36/7=5.14 49/7=7 So the ratio you'd get on C is 5.14:7, which is not what you are looking for. Try to find a number to divide answer D by to get your ratio you are looking for and show your work.

OpenStudy (cpinder):

The answer is D. If you simplify 48:56 by dividing both numbers by 8, you will get your ratio of 6:7.